OpenStudy (loser66):

Statistics problem Let T1,T2 be random times for a company to complete 2 steps in a certain process. Say T1 and T2 are measured in days and they have the joint pdf that is uniform over the space 110)?? Please, help

OpenStudy (perl):

this is a double integral

OpenStudy (loser66):

\(f_X(x) = \dfrac{1}{10-1}=\dfrac{1}{9}\) \(f_Y(y)=\dfrac{1}{6-2}=\dfrac{1}{4}\)

OpenStudy (loser66):

But I don't know how to use t1+2t2 < 14

OpenStudy (perl):

lets first draw the space

OpenStudy (loser66):

but it is wrong, the answer is 1/10. Please, check

OpenStudy (loser66):

@ganeshie8 help me, please

ganeshie8 (ganeshie8):

im getting below region with the given inequalities http://prntscr.com/58yox0

ganeshie8 (ganeshie8):

vertical : t1 horizontal : t2

OpenStudy (loser66):

You mean the bigger one is the area we need to find?

ganeshie8 (ganeshie8):

take the ratio of small to big

ganeshie8 (ganeshie8):

http://prntscr.com/58yr6q

OpenStudy (loser66):

Let me redo it. :)

ganeshie8 (ganeshie8):

okike

ganeshie8 (ganeshie8):

use desmos maybe https://www.desmos.com/calculator

OpenStudy (loser66):

I got it. I know what is my mistake, thanks a lot